# Uni-MuMER: Unified Multi-Task Fine-Tuning of Vision-Language Model for Handwritten Mathematical Expression Recognition

### v2

@@ Abstract
Handwritten Mathematical Expression Recognition (HMER) remains a persistent challenge in Optical Character Recognition (OCR) due to the inherent freedom of symbol layout and variability in handwriting styles. Prior methods have faced performance bottlenecks, proposing isolated architectural modifications that are difficult to integrate coherently into a unified framework. Meanwhile, recent advances in pretrained vision-language models (VLMs) have demonstrated strong cross-task generalization, offering a promising foundation for developing unified solutions. In this paper, we introduce Uni-MuMER, which fully fine-tunes a VLM for the HMER task without modifying its architecture, effectively injecting domain-specific knowledge into a generalist framework. Our method integrates three data-driven tasks: Tree-Aware Chain-of-Thought (Tree-CoT) for structured spatial reasoning, Error-Driven Learning (EDL) for reducing confusion among visually similar characters, and Symbol Counting (SC) for improving recognition consistency in long expressions. Experiments on the CROHME and HME100K datasets show that Uni-MuMER achieves new state-of-the-art performance, surpassing the best lightweight specialized model SSAN by 16.31% and the top-performing VLM Gemini2.5-flash by 24.42% in the zero-shot setting. Our datasets, models, and code are open-sourced at:   https://github.com/BFlameSwift/Uni-MuMER
footnotetext: Equal contribution, Corresponding authors.
@@ 1 Introduction
Handwritten Mathematical Expression Recognition (HMER) seeks to translate handwritten expressions into machine-readable markup, supporting document understanding and digital preservation of scientific material. Unlike standard Optical Character Recognition (OCR), HMER involves parsing complex 2D structures [16], ambiguous symbols, and the inherent freedom of symbol layout and variability in handwriting styles [68], requiring not just recognizing individual symbols but also layout reasoning and parsing their complex spatial relationships. Recent approaches have primarily leveraged RNNs [23, 14] or Transformers [49] due to their powerful sequential modeling capabilities. Additional modules, such as tree-structured decoders [64, 53, 70] and relative-position-aware mechanisms [19, 62, 54, 55], have been introduced to further enhance performance. These dominant research directions in HMER have focused on incorporating prior human knowledge into model architectures, particularly through manually designed structural modules or attention mechanisms [63, 65, 66, 67, 30, 19, 3].
@@ 3.1 Vanilla HMER
Vanilla HMER  As is shown at the top of Fig. 2, Vanilla HMER involves providing an image of a mathematical expression alongside a textual instruction, prompting the model to directly output the corresponding   formatted expression. Traditional lightweight specialized models took images alone as input and generated the expressions, whereas VLMs now exhibit strong generalization capabilities across related tasks. We also utilize models fine-tuned specifically on the Vanilla HMER task as baselines for subsequent comparison. Through Vanilla HMER, the model acquires essential recognition capabilities and ensures accurate and structural outputs.
@@ 3.2 Tree-Aware Chain-of-Thought
Tree-CoT  The construction procedure of Tree-CoT is illustrated in Fig. 3, we first parse the input expression to derive its Abstract Syntax Tree (AST), explicitly capturing hierarchical and spatial relationships among symbols. We subsequently perform a depth-first traversal (DFS) of the AST to sequentially linearize the tree structure. To encode this structured representation, we introduce a specific serialization scheme, using tab-based indentation to reflect tree depth and newline separation to distinguish individual nodes clearly, and the raw text is in Fig. 2. Each serialized line explicitly contains the symbol label and its spatial relation to its immediate parent node, resulting in a coherent textual representation of the AST.
@@ 3.3 Error-Driven Learning
Error-Driven Learning (EDL) leverages a learning-from-mistakes paradigm to enhance model accuracy. Its input consists of an error corpus generated by the model itself, structured into two distinct subtasks: error detection and error correction. Inspired by self-improvement strategies and previous work indicating the suitability of language models for correcting HMER errors [8], we integrate these subtasks explicitly. The subsequent sections detail the generation of the error corpus, and the definitions of error detection and correction subtasks.
@@
Error Corpus Generation  As is shown in the Fig. 4. Initially, the complete dataset (e.g., CROHME) is randomly partitioned into multiple distinct subsets denoted as $F_{1},F_{2},\dots,F_{n}$. Subsequently, we perform cross-validation training, where each VLM is trained on $n-1$ folds and evaluated on the remaining held-out fold $F_{i}$. During this, multiple candidate predictions are generated through multiple sampling for each input image to collect potential model outputs. By comparing these candidate predictions to their corresponding ground-truth labels, we explicitly identify erroneous outputs and compile these erroneous predictions into the final error corpus.
@@ 4.1 Experimental Setup
Datasets  The CROHME series [38, 39, 37] is the most widely used dataset for HMER. It comprises 8836 training samples, and three test sets contain 986, 1147, and 1199 images for the CROHME 2014, 2016, 2019. Recently, the CROHME 2023 [56] iteration provided an expanded training set of 10979 images and a test set of 2300 images. HME100K  [59] is a large-scale, real-world dataset. It consists of 74502 training and 24607 testing images, encompassing various writing styles and conditions. MathWriting  [18], released by Google Research in 2024, is the largest HME corpus to date with 230K human-written and 400K synthetically-generated expressions. Finally, Im2Latexv2[45] builds upon the printed expression dataset Im2Latex-100k[12] and enhances it with improved  normalization and 61 diverse rendering styles for more realistic expression images.
@@
Training  We fine-tune the Qwen2.5-VL-3B model on all datasets with three data-driven tasks simultaneously for a single epoch. Further details can be found in the Appendix.
@@ 4.2 Main Results
Exp\,$\uparrow$ @CDM\,$\uparrow$ Exp\,$\uparrow$ @CDM\,$\uparrow$ Exp\,$\uparrow$ @CDM\,$\uparrow$ Exp\,$\uparrow$ @CDM\,$\uparrow$
@@
DenseWAP[63] 50.1 - 47.5 - 52.96 - 50.19 -
BTTR[67] 53.96 - 52.31 - 52.96 - 53.08 -
ABM[3] 56.85 - 52.92 - 53.96 - 54.58 -
CAN-ABM[30] 57.26 - 56.15 - 55.96 - 56.46 -
@@
SSANa[62] 62.58 - 62.51 - 65.30 - 63.43 -
VLPG[20] 60.41 - 60.51 - 62.34 - 61.09 -
HiE-VLa[21] 73.30 - 70.70 - 69.3 - 71.13 -
Vanilla (baseline) 72.62 75.66 64.26 67.83 69.14 71.98 68.67 71.82
Uni-MuMER 75.36+2.74 79.11+3.45 70.79+6.53 75.24+7.41 73.73+4.59 77.23+5.25 73.29+4.62 77.19+5.37
Uni-MuMER 82.05+9.43 85.09+9.43 77.94+13.68 81.08+13.25 79.23+10.09 82.40+10.42 79.74+11.07 82.86+11.04
Tab. 1 displays the strong performance of our Uni-MuMER and previous methods on the CROHME dataset. As CDM is a recently proposed metric, we reproduced all reported results involving CDM. For a fair comparison, the Vanilla(baseline) reported in Tab. 1 and Tab. 2 was trained exclusively on either CROHME or HME100K dataset. The Uni-MuMER variant incorporated training samples constructed from the corresponding images for three tasks. Further utilizing external data, UniMuMER was trained on approximately 1.6M samples, comprising data from 392K images across the three tasks and all datasets.
Comparison with Previous SOTA Methods  Data augmentation has traditionally improved performance in lightweight models, yet inherent limitations still restrict overall effectiveness. Recently, VLPG [20] and HiE-VL [21], have integrated VLM into HMER. HiE-VL enhances performance through non-end-to-end training, architectural refinements, and extra data. However, both Uni-MuMER and Uni-MuMER (use external data) maintain a notable performance advantage over these VLM-based methods. Notably, we also trained CoMERa [66] using the same external data (392K images) and observed a significant performance degradation. This finding indicates that the inherent limitations of lightweight models hinder their ability to effectively utilize large and diverse datasets.
@@
Uni-MuMER 71.93+1.78 73.67+1.87
Incremental Improvement  Building upon the Qwen2.5-VL-3B VLM base, our Vanilla(baseline) model already surpasses most published SOTA methods. After integrating three data-driven tasks, Tree-CoT, EDL, and SC, our method, Uni-MuMER further improves both ExpRate and ExpRate@CDM, as shown in Tab. 1. For a fair comparison, both the Vanilla (baseline) and Uni-MuMER are trained solely on the CROHME training set, without using any external data. Introducting Upon incorporating external data along with these tasks(1.6M samples), our final model, Uni-MuMER, achieves even greater performance, notably reaching an ExpRate of 79.74% and ExpRate@CDM of 82.86% on the CROHME Average. Both Uni-MuMER and Uni-MuMER significantly outperform both the Vanilla (baseline) model and prior methods, underscoring the vital role of data-driven tasks and diverse external data in enhancing model accuracy.
Performance on HME100K  To focus the evaluation on mathematical expression recognition, following [17], we exclude all test instances containing CJK characters. Tab. 2 shows that open-source and closed-source VLMs exhibit modest performance, which is due to the challenging real-world conditions in HME100K dataset. In contrast, several prior lightweight specialized methods exclusively trained on HME100K, such as TAMER, achieve strong performance. Although HiE-VL utilizes external training data, it achieves limited results on HME100K, even underperforming compared to lightweight specialized models. Our Uni-MuMER achieves SOTA performance. Additionally, by using external data, Uni-MuMER further enhances its performance, clearly surpassing HiE-VL.
@@ 5.1 Ablation Study
(a) Model performance relative to structural complexity. Tree-CoT task notably reduces accuracy loss as structural complexity increases.
@@
5.2 Data Diversity Effects
@@
Necessity of Domain-Specific Training  We observe that large differences in handwriting style across datasets severely limit cross-domain generalization. A model trained only on CROHME data excels on similar CROHME-like benchmarks but struggles on markedly different handwriting, such as that in the diverse MathWriting dataset, unless explicitly exposed to these domains during training.
@@
Out-of-Domain Generalization  Accuracy on printed expression images gradually improves with increased handwritten training data, reflecting partial transfer of structural knowledge for expression recognition. However, performance on printed expressions remains limited with exclusively handwritten training data exhibiting a marked increase only upon incorporating the Im2LaTeXv2 dataset. This indicates that generalized improvements across fundamentally different domains necessitate explicit representation in the training process.
@@ 6 Discussion
For Small Models: Larger Model, Better Performance, Faster Speed  Previous HMER models were lightweight and relied on task-specific architectural designs (e.g., tree-structured modules). Uni-MuMER builds upon both these task-specific insights and current VLM, taking a significant step forward with a unified and more powerful model. Benefiting from recent advances in efficient inference frameworks, Uni-MuMER not only achieves substantially improved performance but also delivers faster inference compared to smaller models. Further analysis is provided in the Appendix.
For VLM: A Chain-of-Thought Perspective on Expression Recognition  Unlike direct expression recognition, Uni-MuMER adopts a chain-of-thought perspective by formulating three tasks: expression tree construction, error detection and correction, and symbol counting. These tasks guide model step by step toward the final prediction, enhancing both interpretability and generalization. This approach offers a novel direction for advancing expression recognition with current VLM.
Future Direction: Self-Correction for Expression Recognition  Uni-MuMER provides a new perspective for enabling self-correction during inference in expression recognition. By integrating the Tree-CoT mechanism with an explicit error correction task, the model can exhibit self-corrective behavior within CoT process. Moreover, incorporating colloquial, human-like expressions (e.g., "wait, there’s a left parenthesis earlier—this is more likely a right parenthesis, not ’2’") can further induce R1-like spontaneous self-correction[11], enhancing the upper limit of the model’s performance.
Limitations  Despite the strong results, our method has several limitations. Due to computational resource constraints, a 400k synthetic subset of the MathWriting dataset was unused, potentially limiting the diversity benefits. Besides, we don’t evaluate Uni-MuMER on general VLM benchmarks, such as OCR. Integrating our method with general VLM fine-tuning datasets and assessing its performance on broader benchmarks remains an important area for future analysis.

### v3

@@ Abstract
Handwritten Mathematical Expression Recognition (HMER) remains a persistent challenge in Optical Character Recognition (OCR) due to the inherent freedom of symbol layouts and variability in handwriting styles. Prior methods have faced performance bottlenecks by proposing isolated architectural modifications, making them difficult to integrate coherently into a unified framework. Meanwhile, recent advances in pretrained vision-language models (VLMs) have demonstrated strong cross-task generalization, offering a promising foundation for developing unified solutions. In this paper, we introduce Uni-MuMER, which fully fine-tunes a VLM for the HMER task without modifying its architecture, effectively injecting domain-specific knowledge into a generalist framework. Our method integrates three data-driven tasks: Tree-Aware Chain-of-Thought (Tree-CoT) for structured spatial reasoning, Error-Driven Learning (EDL) for reducing confusion among visually similar characters, and Symbol Counting (SC) for improving recognition consistency in long expressions. Experiments on the CROHME and HME100K datasets show that Uni-MuMER achieves super state-of-the-art performance, outperforming the best lightweight specialized model SSAN by 16.31% and the top-performing VLM Gemini2.5-flash by 24.42% under zero-shot setting. Our datasets, models, and code are open-sourced at:   https://github.com/BFlameSwift/Uni-MuMER
[…]
@@ 1 Introduction
Handwritten Mathematical Expression Recognition (HMER) seeks to translate handwritten expressions into machine-readable markup, supporting document understanding and digital preservation of scientific material. Unlike standard Optical Character Recognition (OCR), HMER involves parsing complex 2D structures [15], ambiguous symbols, and the inherent freedom of symbol layout and variability in handwriting styles [68], requiring not just recognizing individual symbols but also layout reasoning and parsing their complex spatial relationships. Recent approaches have primarily leveraged RNNs [23] or Transformers [48] due to their powerful sequential modeling capabilities. Additional modules, such as tree-structured decoders [64, 52, 70] and relative-position-aware mechanisms [18, 62, 53, 54], have been introduced to further enhance performance. These dominant research directions in HMER have focused on incorporating prior human knowledge into model architectures, particularly through manually designed structural modules or attention mechanisms [63, 65, 66, 67, 30, 18, 4].
@@ 3.1 Vanilla HMER
Vanilla HMER  As is shown at the top of Fig. 2, Vanilla HMER involves providing an image of a mathematical expression alongside a textual instruction, prompting the model to directly output the corresponding formatted expression. Traditional lightweight specialized models took images alone as input and generated the expressions, whereas VLMs now exhibit strong generalization capabilities across related tasks. We also utilize models fine-tuned specifically on the Vanilla HMER task as baselines for subsequent comparison. Through Vanilla HMER, the model acquires essential recognition capabilities and ensures accurate and structured outputs.
@@ 3.2 Tree-Aware Chain-of-Thought
Tree-CoT  The construction procedure of Tree-CoT is illustrated in Fig. 3. We first parse the input expression to derive its Abstract Syntax Tree (AST), explicitly capturing hierarchical and spatial relationships among symbols. We subsequently perform a depth-first traversal (DFS) of the AST to sequentially linearize the tree structure. To encode this structured representation, we introduce a specific serialization scheme, using tab-based indentation to reflect tree depth and newline separation to distinguish individual nodes clearly, and the raw text is in Fig. 2. Each serialized line explicitly contains the symbol label and its spatial relation to its immediate parent node, resulting in a coherent textual representation of the AST.
@@ 3.3 Error-Driven Learning
Error-Driven Learning (EDL) leverages a learning-from-mistakes paradigm to enhance model accuracy. Its input consists of an error corpus generated by the model itself, structured into two distinct subtasks: error detection and error correction. Inspired by self-improvement strategies and previous work indicating the suitability of language models for correcting HMER errors [9], we integrate these subtasks explicitly. The subsequent sections detail the generation of the error corpus and the definitions of error detection and correction subtasks.
@@
Error Corpus Generation  As is shown in the Fig. 4. Initially, the complete dataset (e.g., CROHME) is randomly partitioned into multiple distinct subsets denoted as $F_{1},F_{2},\dots,F_{n}$. Subsequently, we perform cross-validation training, where each VLM is trained on n1n-1 folds and evaluated on the remaining held-out fold $F_{i}$. During this, multiple candidate predictions are generated through multiple sampling for each input image to collect potential model outputs. By comparing these candidate predictions to their corresponding ground-truth labels, we explicitly identify erroneous outputs and compile these erroneous predictions into the final error corpus.
@@ 4.1 Experimental Setup
Datasets  The CROHME series [37, 38, 39] is the most widely used dataset for HMER. It comprises 8836 training samples, and three test sets contain 986, 1147, and 1199 images for the CROHME 2014, 2016, 2019. Recently, the CROHME 2023 [56] iteration provided an expanded training set of 10979 images and a test set of 2300 images. HME100K  [59] is a large-scale, real-world dataset. It consists of 74502 training and 24607 testing images, encompassing various writing styles and conditions. MathWriting  [17], released by Google Research in 2024, is the largest HME corpus to date with 230K human-written and 400K synthetically-generated expressions. Finally, Im2Latexv2[44] builds upon the printed expression dataset Im2Latex-100k[12] and enhances it with improved normalization and 59 diverse rendering styles for more realistic expression images.
@@
Training  We fine-tune the Qwen2.5-VL-3B model on all datasets with three data-driven tasks simultaneously for a single epoch. Further details can be found in the Appendix B.4.
@@ 4.2 Main Results
$uparrow Exp\,\uparrow @CDM\,\uparrow$ $uparrow Exp\,\uparrow @CDM\,\uparrow$
Closed-Source Models
@@
Specialized Models
DenseWAP[63] 50.1 47.5
BTTR[67] 53.96 52.31 52.96 53.08
ABM[4] 56.85 52.92 53.96 54.58
CAN-ABM[30] 57.26 56.15 55.96 56.46
@@
SSANa[62] 62.58 62.51 65.30 63.43
UniMERNeta 67.4 68.4 65.4 67.07
VLPG[20] 60.41 60.51 62.34 61.09
HiE-VLa[21] 73.30 70.70 69.3 71.13
Vanilla (baseline) 72.62 75.66 64.26 67.83 69.14 71.98 68.64 71.82
Uni-MuMER 75.36+2.74 79.11+3.45 70.79+6.53 75.24+7.41 73.73+4.59 77.23+5.25 73.29+4.65 77.19+5.37
Uni-MuMER 82.05+9.43 85.09+9.43 77.94+13.68 81.08+13.25 79.23+10.09 82.40+10.42 79.74+11.10 82.86+11.04
Tab. 1 displays the strong performance of our Uni-MuMER and previous methods on the CROHME dataset. As CDM is a recently proposed metric, we reproduced all reported results involving CDM. For a fair comparison, the Vanilla (baseline) reported in Tab. 1 and Tab. 2 was trained exclusively on either CROHME or HME100K dataset. The Uni-MuMER variant incorporated training samples constructed from the corresponding images for three tasks. Further utilizing external data, UniMuMER was trained on approximately 1.6M samples, comprising data from 386K images across the three tasks and all datasets. Due to space limitations, additional results are presented in Appendix D.
Comparison with Previous SOTA Methods  Data augmentation has traditionally improved performance in lightweight models, yet inherent limitations still restrict overall effectiveness. Recently, VLPG [20] and HiE-VL [21], have integrated VLM into HMER. HiE-VL enhances performance through non-end-to-end training, architectural refinements, and extra data. However, both Uni-MuMER and Uni-MuMER (use external data) maintain a notable performance advantage over these VLM-based methods. Notably, we also trained CoMERa [66] using the same external data (386K images) and observed a significant performance degradation. This finding indicates that the inherent limitations of lightweight models hinder their ability to effectively utilize large and diverse datasets.
@@
Uni-MuMER 72.66+2.51 74.30+2.50
Incremental Gains  Building upon the Qwen2.5-VL-3B VLM base, our Vanilla (baseline) model already surpasses most published SOTA methods. After integrating three data-driven tasks, Tree-CoT, EDL, and SC, our method, Uni-MuMER further improves both ExpRate and ExpRate@CDM, as shown in Tab. 1. For a fair comparison, both the Vanilla (baseline) and Uni-MuMER are trained solely on the CROHME training set, without using any external data. Introducting Upon incorporating external data along with these tasks (1.6M samples), our final model, Uni-MuMER, achieves even greater performance, notably reaching an ExpRate of 79.74% and ExpRate@CDM of 82.86% on the CROHME Average. Both Uni-MuMER and Uni-MuMER significantly outperform both the Vanilla (baseline) model and prior methods, underscoring the vital role of data-driven tasks and diverse external data in enhancing model accuracy.
Performance on HME100K  To focus the evaluation on mathematical expression recognition, following [16], we exclude all test instances containing CJK characters. Detailed impacts are discussed in Appendix C.5. Tab. 2 shows that open-source and closed-source VLMs exhibit modest performance, which is due to the challenging real-world conditions in HME100K dataset. In contrast, several prior lightweight specialized methods exclusively trained on HME100K, such as TAMER, achieve strong performance. Although HiE-VL utilizes external training data, it achieves limited results on HME100K, even underperforming compared to prior models. Our Uni-MuMER achieves SOTA performance. Additionally, by using external data, Uni-MuMER further enhances its performance, clearly surpassing HiE-VL.
@@ 5.1 Ablation Study
(a) Model performance relative to structural complexity. The Tree-CoT task notably reduces accuracy loss as structural complexity increases.
@@
5.2 Effects of Data Diversity
@@
Out-of-Domain Generalization  Accuracy on printed expression images gradually improves with increased handwritten training data, reflecting partial transfer of structural knowledge for expression recognition. Nevertheless, its accuracy improves significantly after adding Im2LaTeXv2, reflecting the value of in‑domain printed data for that target.
Leave-One-Out Ablation  To further investigate the out-of-domain generalization performance, we conducted an additional “Leave-One-Out” experiment, training Uni-MuMER on N-1 datasets and evaluating it on the held-out N-th dataset. The results below present the expression recognition accuracy in each scenario. For comparison, we include the baseline CoMER and Doubao-1.5-pro under zero-shot settings.
Table 5: Leave-one-out ablation of training datasets. (ExpRate@CDM%) We train Uni-MuMER in six configurations: using all data (all five sources), and using all-minus-one for each major dataset. Results are reported on CROHME-Average, CROHME 2023, HME100K, MathWriting, and Im2LaTeXv2. For reference, we also show a specialized model (CoMER) trained with and without CROHME, and a closed-source VLM (Doubao-1.5-pro) for reference.
Model CROHME-Avg CROHME 2023 HME100K MathWriting Im2LaTeXv2
CoMER (w/o CROHME) 45.99 58.28 39.49 23.23 40.39
CoMER 52.29 59.91 44.63 28.45 53.37
Doubao-1.5-pro 65.77 58.18 55.08 26.34 27.66
Uni-MuMER (w/o CROHME) 78.22 72.19 74.47 68.08 89.18
Uni-MuMER (w/o CROHME-2023) 82.31 71.66 74.68 68.85 89.77
Uni-MuMER (w/o HME100K) 82.30 72.01 52.38 68.94 91.19
Uni-MuMER (w/o MathWriting) 79.58 73.51 74.29 32.55 93.31
Uni-MuMER (w/o Im2LaTeXv2) 83.20 73.33 74.51 68.93 75.17
Uni-MuMER 82.89 74.42 74.30 69.11 93.29
The leave-one-out evaluation confirms that our unified model can generalize to an unseen dataset reasonably well. When the unseen domain is very different, performance drops more markedly, which highlights an area for improvement. HME100K for real-life, low-quality images, MathWriting for densely structured expression, and Im2LaTeXv2 for printed expression images, each of which represents a markedly different and challenging domain.
Crucially, after introducing even modest amounts of domain-relevant data, our unified approach easily achieves top-tier performance. Thus, our results indicate that, compared to existing specialized models and general VLM baselines, our approach provides superior generalization with minimal domain-specific data requirements, highlighting a clear advantage and direction for future improvement.
5.3 Mixing General‑Purpose VLM Data
From the perspective of general‑purpose LLMs and VLMs, mixing data from different domains can yield substantial performance gains. Although our Uni-MuMER addresses previous limitations by introducing a unified solution, achieving super SOTA performance, it’s necessary to validate whether Uni‑MuMER benefits from general domain data beyond HMER‑only training and how Uni-MuMER-Data affects the model’s general capabilities. To this end, we introduce Uni-MuMER-LLAVA, fine-tuned using an equal 1:1 mixture of HMER and LLaVA-OneVision data, and compare its performance with Uni-MuMER across benchmarks assessing HMER-specific and general capabilities.
Table 6: Comparison among Uni-MuMER, Uni-MuMER-LLAVA and Qwen2.5-VL on multi-modal reasoning benchmarks. Results are reported on MMMU_val, Math-Vision, and MathVista_testmini, covering general multimodal understanding and mathematical reasoning tasks.
Model MMMU_val Math-Vision MathVista_testmini
Uni-MuMER 47.89 24.01 47.8
Uni-MuMER-LLAVA 48.67 24.34 51.1
Qwen2.5-VL-3B-Instruct 52.78 21.38 47.9
Table 7: Cross-benchmark comparison between Uni-MuMER and Uni-MuMER-LLAVA across benchmarks. Results are reported on CROHME-Avg, CROHME-2023, HME100K, MathWriting, and Im2LaTeXv2.
Model CROHME-Avg CROHME-2023 HME100K MathWriting Im2LaTeXv2
Uni-MuMER 82.89 74.42 74.30 69.11 93.29
Uni-MuMER-LLAVA 81.99 69.10 73.48 70.48 93.44
As shown in Tab. 6 and Tab. 7, Uni-MuMER exhibits strong generalization capabilities despite being fine-tuned exclusively on HMER-specific data, achieving performance comparable to Qwen2.5VL-3B. Notably, Uni-MuMER-LLAVA further enhances overall performance, surpassing Uni-MuMER on general tasks like MMMU and Math, while incurring marginal performance decreases on specific HMER test sets. It’s promising to integrate our method and Uni-MuMER-Data to empower VLMs.
@@ 6 Discussion
For Small Models: Larger Model, Better Performance, Faster Speed  Previous HMER models were lightweight and relied on task-specific architectural designs (e.g., tree-structured modules). Uni-MuMER builds upon both these task-specific insights and current VLM, taking a significant step forward with a unified and more powerful model. Benefiting from recent advances in efficient inference frameworks, Uni-MuMER not only achieves substantially improved performance but also delivers faster inference compared to smaller models. Further detail is provided in Appendix C.2.
For VLM: A Chain-of-Thought Perspective on Expression Recognition  Unlike direct expression recognition, Uni-MuMER adopts a chain-of-thought perspective by formulating three tasks: expression tree construction, error detection and correction, and symbol counting. These tasks guide the model step by step toward the final prediction, enhancing both interpretability and generalization. This approach offers a novel direction for advancing expression recognition with current VLM.
Future Direction: Self-Correction for Expression Recognition  Uni-MuMER provides a new perspective for enabling self-correction during inference in expression recognition. By integrating the Tree-CoT mechanism with an explicit error correction task, the model can exhibit self-corrective behavior within CoT process. Moreover, incorporating colloquial, human-like expressions (e.g., ”wait, there’s a left parenthesis earlier—this is more likely a right parenthesis, not ’2’”) can further induce R1-like spontaneous self-correction [19], enhancing the upper limit of the model’s performance.
Limitations  Despite the strong results, our method has several limitations. Due to computational resource constraints, a 400k synthetic subset of the MathWriting dataset was unused, potentially limiting the diversity benefits. Additionally, exploring optimal task mixture ratios within Uni-MuMER and optimal domain mixture ratios between HMER-specific data and general-domain data is crucial; however, these experiments were omitted due to significant resource constraints. Investigating these ratios could yield further enhance performance and generalizability.